Reconsider problem 1 but suppose that the sample is selected with replacement. A small pond contains 30 fish, 10 of which have been tagged. Suppose that a fisherman’s catches four fish by catching a fish, noting whether it is tagged, releasing it, and repeating until four fish have been caught. (Assume that these four fish form a random sample of size four selected with replacement.) Let X denote the number of tagged fish that are caught.

a) What is the name of the distribution of X and what is the p.m.f. of X for this example? Provide an expression for pX(x).

b) Find the probability that the fisherman’s catch will contain at most 3 tagged fish. c) Use the appropriate expressions from section 7.9 to compute expected value of X and the variance of X.
Answer #1

(a)

Since out of 30 fishes, 10 are tagged so the probability that a fish is tagged is

p = P(tagged) = 10/30 = 1/3 = 0.3333

Since fish are drawn with replacement so each time probability of getting tagged fish remain same.

Here X has binomial distribution with parameter n= 4 and p = 0.3333.

The expression of pmf is

p_{X}(x)=\binom{4}{x}(0.3333)^{x}(1-0.3333)^{30-x},x=0,1,2,3,4

Following table shows the pmf:

X p(x)
0 0.19757
1 0.39508
2 0.29627
3 0.09874
4 0.01234

(b)

P(X\leq 3)=1-P(X>3)=1-p_{X}(4)=0.98766

(c)

Following table shows the calculations:

X p(x) x*p(x) x^2*p(x)
0 0.19757 0 0
1 0.39508 0.39508 0.39508
2 0.29627 0.59254 1.18508
3 0.09874 0.29622 0.88866
4 0.01234 0.04936 0.19744
Total 1.3332 2.66626

The mean is

\mu=E(X)=\sum xp_{X}(x)=1.3332

The variance is

\sigma^{2}=E(X^{2})-[E(X)]^{2}=\sum x^{2}p_{X}(x)-\left [\sum xp_{X}(x) \right ]^{2}=0.8888

Another way to find mean and variance:

\mu=np=4\cdot 0.3333=1.3332

\sigma^{2}=np(1-p)=4\cdot 0.3333\cdot (1-0.3333)=0.88884444\approx 0.8888